10a + 2b = 15 and 5a = 4b, what is the value of a + b?
Answer by oberobic(2304)

10a + 2b = 15
5a - 4b = 0
.
Multiply 2nd equation by 2 and then subtract
.
10a + 2b = 15
10a - 8b = 0
10b = 15
b = 1.5
.
5a = 4b
5a = 4*1.5 = 6
a = 1.2
.
Checking...does 10a + 2b = 15?
.
10a = 10*1.2= 12
2b= 2*1.5 = 3
12+3 = 15
Check.
.
Answer:
a + b = 1.2+1.5 = 2.7
.
Done.